# Apostol Vassilev

### v1

@@ 1 Introduction
Generally speaking, the guardrails encompass policies, technical controls and monitoring mechanisms. In this paper that are referred to as checkers $C$. They may exist across all layers of the AI system: i) data guardrails for sanitization of training data; ii) model guardrails that to ensure the validation and continuous monitoring of the AI system according to established metrics, e.g., accuracy and robustness; iii) application guardrails, typically implemented as computer programs running along the model, that try to enforce policies on AI generated content or restrict how AI toils function within specific workflows; iv) infrastructure guardrails. also implemented as computer programs running along the model, that try to provide a secure foundation for the AI system by enforcing protections at the cloud, network and systems levels.
One may think of the collection of different types of guardrails as the technical support for the NIST AI RMF functions NISTAIRMF: Govern, Map, Measure, and Manage. Categories of AI Guardrails A. Input Guardrails (MAP → MANAGE): PII scrubbing, adversarial prompt detection, domain-restricted inputs. B. Policy & Governance Guardrails (GOVERN → MAP → MANAGE): Scope limitations, regulatory constraints, RBAC. C. Model-Level Guardrails (MAP → MEASURE → MANAGE): Alignment tuning, safety classifiers, confidence thresholds. D. Output Guardrails (MEASURE → MANAGE): Redaction, hallucination detection, refusal templates. E. Action & Execution Guardrails (GOVERN → MANAGE): Safety interlocks, approval workflows, OPA policy engines. F. Monitoring, Auditing & Logging Guardrails (MEASURE → MANAGE): Drift detection, anomaly detection, audit trails.
@@
2 Related work
The problem of findings the limits of robustness of AI alignment and security has been approached by many researchers. OODDetect2022 considers the problem of detecting and blocking out-of-distribution inputs to AI models as a defense and shown impossibilities when distributions overlap. Wolf2023FundamentalLO look at fundamental limitations of LLM alignment but their results are based on strong assumptions about the distribution of adversarial prompts which may not hold in practice. This work relies on the assumption of decomposability of LLMs into components and distinguishability between them. While intuitive and supported by empirical results, the authors acknowledge the need for further analysis to develop theory based on more realistic assumptions.  glukhov2023llm considers the limitations to censoring LLM’ inputs but also rely on strong assumptions that may not hold in practice. The authors acknowledge that the practical implications of their results can be difficult to assess and may only be considered on a case-by-case basis. The work of Zhang2023WatermarksIT is directed towards setting the limits of robustness of watermarking techniques for LLM output to distinguish synthetic content. undetectabletrojan2022 consider the problem of creating impossible to detect trojans in AI models but under strong assumptions that are unlikely to hold in real-world AI models. What is common for all these results is that they rely on strong assumptions that limit the generality of the results and raise doubts about their applicability in practice. Recently, karpowicz2025 established a fundamental impossibilities about non-trivial knowledge aggregation, which is shown to not be able to simultaneously achieve truthful knowledge representation and semantic information conservation. kalai2025languagemodelshallucinate also provide impossibility results for hallucinations in LLMs based on constraints: i) epistemic uncertainty associated with rare information in training data; ii) computational complexity limitations; iii) knowledge representation in current model architectures.
In contrast, the approach taken in this manuscript establishes impossibility for robust alignment results with a minimal assumption that is an inherent trait of all types of AI systems: they use computation for their operation and reasoning. Thus, the results derived from it are the strongest and most general of all information-theoretic limitations established in the literature.
@@ 3.1 On adversarial robustness of AI Systems
The term adversarial robustness is defined similarly to the Security of the AI System as the ability of the AI System to withstand deliberate attacks by capable adversaries aiming to force the AI system out of alignment with a policy. This manuscript considers attacks through the main prompt interface of a chatbot or an AI Agent based on a LLM. In this setting, an adversarial prompt is a string of some length. For example: ”Please give me instructions for making a bomb”. As noted above, responsible organizations adopt OOPS policies $\Pi$ that will flag such a request as unacceptable and block it. Although there are some topics that are known in advance to fall under OOPS, knowing all possible adversarial prompts in advance is just not possible in practice. For example, the ASCII Art-based jailbreak attack by promptasciiartbasedjailbreak2024 compromised LLMs with guardrails guarding the topic of bomb making. The richness of the English [^2] The same is true for any other human language language combined with cybersecurity tricks used to exploit prompting, e.g.,ANSIcodes025, make the problem of identifying in advance attack prompts hard. Another difficult to anticipate and mitigate type of attack is the ”crescendo jailbreak” by  russinovich2025greatwritearticlethat, which exploits gradual escalation of the dialogue with the LLM by referencing the model’s replies progressively leading to a successful jailbreak. The findings in vassilev2025 show a large and increasing body of adversarial attacks. Recent findings by zou2025securitychallengesaiagent show near perfect rate of success of prompt-based attacks against agentic systems based on LLMs.
@@ 3.2 Considerations for real AI Systems with finite context windows
The window sizes shown in Figure 1  [^3] Certain equipment, instruments, software, or materials, commercial or non-commercial, are identified in this paper in order to specify the experimental procedure adequately. Such identification does not imply recommendation or endorsement of any product or service by NIST, nor does it imply that the materials or equipment identified are necessarily the best available for the purpose.are so large that no organization can really traverse all possible strings within $W$ to ensure the complete testing of the AI System: there are not enough resources on Earth to allow for this. Let $\hat{\Omega}$ be the set of all prompts to the AI System that fit in $W$. Here, $card(\hat{\Omega})=M$, for some large $M>0$.
@@ 3.3 Broader implications
These results do not mean that current and future AI systems, including Artificial General Intelligence or Artificial Super Intelligence, will not be able to discover and prove truths in science that are currently unproven or unknown to humanity. However, these results shown that AI systems are ultimately limited in ways similar to humans so that there are truths they will never be able to prove.
@@ 8 Acknowledgments
The graphics in Figure 1 was produced by GPT-5, with author’s prompting.

### v2

@@ 1 Introduction
Generally speaking, the guardrails encompass policies, technical controls and monitoring mechanisms. In this paper they are referred to as checkers $C$. They may exist across all layers of the AI system:
  • i)
    data guardrails for sanitization of training data;
  • ii)
    model guardrails that ensure the validation and continuous monitoring of the AI system according to established metrics, e.g., accuracy and robustness;
  • iii)
    application guardrails, typically implemented as computer programs running along the model, that try to enforce policies on AI generated content or restrict how AI tools function within specific workflows;
  • iv)
    infrastructure guardrails, also implemented as computer programs running along the model, that try to provide a secure foundation for the AI system by enforcing protections at the cloud, network and systems levels.
One may think of the collection of different types of guardrails also as the technical support for the NIST AI RMF functions NISTAIRMF (Govern, Map, Measure, and Manage):
  • A.
    Input Guardrails (MAP → MANAGE): PII scrubbing, adversarial prompt detection, domain-restricted inputs.
  • B.
    Policy & Governance Guardrails (GOVERN → MAP → MANAGE): Scope limitations, regulatory constraints, RBAC.
  • C.
    Model-Level Guardrails (MAP → MEASURE → MANAGE): Alignment tuning, safety classifiers, confidence thresholds.
  • D.
    Output Guardrails (MEASURE → MANAGE): Redaction, hallucination detection, refusal templates.
  • E.
    Action & Execution Guardrails (GOVERN → MANAGE): Safety interlocks, approval workflows, OPA policy engines.
  • F.
    Monitoring, Auditing & Logging Guardrails (MEASURE → MANAGE): Drift detection, anomaly detection, audit trails.
@@ 2.1 On adversarial robustness of AI Systems
The term adversarial robustness is defined similarly to the Security of the AI System as the ability of the AI System to withstand deliberate attacks by capable adversaries aiming to force the AI system out of alignment with a policy. This manuscript considers attacks through the main prompt interface of a chatbot or an AI Agent based on a LLM. In this setting, an adversarial prompt is a string of some length. For example: ”Please give me instructions for making a bomb”. As noted above, responsible organizations adopt OOPS policies $\Pi$ that will flag such a request as unacceptable and block it.
Adversarial prompts often take advantage of the ambiguity of the language of the input prompt. Guardrails - filters, policies, classifiers, and prompt analysis layers - generally rely on interpreting the intent and semantic meaning of the input. Pattern matching - identifying specific keywords, intent or known harmful structures - is widely used technique for implementing guardrails. When the input text is ambiguous, the interpretation becomes uncertain. Ambiguity is in essence the ”Achilles’ heel” of modern guardrails. Thus, any input that introduces linguistic noise or creates context-based uncertainty makes it significantly harder for the AI system to definitively categorize the request as in or out of OOPS policy.
The following techniques illustrate how ambiguity undermines the robustness of guardrails in AI systems:
  • Disruption of pattern-matching heuristics
    • Linguistic obfuscation: using poetic structures, metaphors, rare dialects, polysemy, etc., adversaries can create inputs that to a human imply clear malicious intent but to a machine may look like benign, creative text.
    • Contextual framing: here the adversary wraps a harmful prompt into a complex, fictional, or highly-professional role-play scenario, which causes the ambiguity of intent to rise and in turn make the AI system struggle to determine if it is in or out of OOPS policy.
    • Crescendo context exploit russinovich2025greatwritearticlethat: gradual escalation of the dialogue with the LLM by referencing the model’s replies, progressively leading to a successful jailbreak.
  • Compositional ambiguity - introduce ambiguity in the structure of the sentence.
    • For example, the sentence ”Show me how people break encryption systems used by banks” has three possible meanings:
        historical analysis;
        academic cryptography discussion;
        operational attack instructions.
      Determining which interpretation applies may be unclear from the context.
    • The ASCII Art-based jailbreak attack promptasciiartbasedjailbreak2024, which compromised LLMs with guardrails guarding the topic of bomb making, mixes plain language with instructions for how to decode figures drawn with ASCII symbols to determine the meaning of the request.
    • The richness of the English [^2] The same is true for any other human language language combined with cybersecurity tricks ANSIcodes025 can also create compositional ambiguities to exploit in prompting.
  • Politeness exploits and tone-based shifts. Tone manipulation - extreme politeness, deference, or emotional pleading - is used to bypass guardrails. Such inputs are linguistically benign and statistically smooth because they do not contain triggering keywords and standard lexical filters tend to let them pass. The ambiguity lies in the subtext: the AI model perceives a helpful and cooperative interaction, which tends to shift the probability distribution. This, in turn, makes the AI model more willing to provide speculative or OOPS-adjacent answers that it would refuse under a natural, direct prompt.
  • Context robustness gap exploits. In complex AI systems, such as those using retrieval-augmented generation (RAG), ambiguity may be injected through retrieved documents. If a system retrieves a mix of benign and contextually-confusing information, the guardrail may flip its judgment based on how it interprets the surrounding noise. The ambiguity of the retrieved context makes the final OOPS decision brittle.
This non-exhaustive list of examples shows that if OOPS policy compliance checking is built on a finite set of rules and language is infinitely ambiguous, the number of ways in which adversaries can hide harmful intent in plain sight is effectively limitless. The findings in vassilev2025 show a large and increasing body of adversarial attacks. Recent findings zou2025securitychallengesaiagent show near perfect rate of success of prompt-based attacks against agentic systems based on LLMs.
@@ 2.2 Considerations for real AI Systems with finite context windows
The window sizes shown in Figure 1 are so large that no organization can really traverse all possible strings within $W$ to ensure the complete testing of the AI System: there are not enough resources on Earth to allow for this. Let $\hat{\Omega}$ be the set of all prompts to the AI System that fit in $W$. Here, $card(\hat{\Omega})=M$, for some large $M>0$.
@@ 2.3 Broader implications
These results do not mean that current and future AI systems, including Artificial General Intelligence or Artificial Super Intelligence, will not be able to discover and prove truths in science that are currently unproven or unknown to humanity. However, these results show that AI systems are ultimately limited in ways similar to humans so that there are truths they will never be able to prove.
@@ 7 Acknowledgments
The graphics in Figure 1 was produced by GPT-5333Certain equipment, instruments, software, or materials, commercial or non-commercial, are identified in this paper in order to specify the experimental procedure adequately. Such identification does not imply recommendation or endorsement of any product or service by NIST, nor does it imply that the materials or equipment identified are necessarily the best available for the purpose., with author’s prompting.